# Euclid: A blue galaxy population and a brightest cluster galaxy in the making in a MaDCoWS2 galaxy cluster candidate††thanks: This paper is published on behalf of the Euclid Consortium

### 07 Mar 2026 v2

@@
We present an example cluster follow-up study with Euclid. Our target, a candidate cluster nicknamed the ‘Puddle’, was initially discovered by the Massive and Distant Clusters of WISE Survey 2 (MaDCoWS2) as a candidate cluster. It was also detected independently as a candidate with both cluster-finding algorithms in Euclid Quick Release 1 (Q1). A Keck MOSFIRE spectrum shows the brightest nucleus is at and is AGN-dominated. We focus our analysis on the galaxy population and the Brightest Cluster Galaxy (BCG), using a combination of Euclid and ancillary photometry. Compared to similar fields, we measure an overdensity of galaxies with in a radius around the BCG. We estimate that % of the completeness-corrected galaxy population is red, which is consistent with some clusters at but lower than others. Euclid imaging reveals that six or seven galaxies appear to be assembling to form the future BCG. Spectral energy distribution (SED) fitting suggests that the merging BCG has a stellar mass of and experienced a short burst of star formation about Myr ago. Its morphology, stellar mass, and star-formation history suggest that the proto-BCG is a more evolved version of the merging core of SPT234956. These systems indicate that multiobject mergers might be a common BCG formation process. Assuming a similar density of mergers in the Euclid Wide Survey, we expect that Euclid will discover approximately 400 assembling BCGs by the end of its mission.
@@
Galaxy clusters sit at the nodes of the cosmic web and are linked by filaments (Bond et al. 1996; Springel et al. 2005; Vogelsberger et al. 2014). As the maxima of the density field in the Universe, galaxy clusters are natural laboratories to test a variety of astrophysical phenomena in different fields: cosmology (e.g. Vikhlinin et al. 2009; Allen et al. 2011; Pierre et al. 2016), plasma and high-energy physics (e.g. McNamara and Nulsen 2007, 2012; Zhuravleva et al. 2014) and galaxy evolution (e.g. von der Linden et al. 2010; Peng et al. 2010b, 2012).
@@
The cessation of star formation in galaxies, usually called ‘quenching,’ depends on two main factors: galaxy mass (Kauffmann et al. 2003) and environment density (Peng et al. 2010b, 2012). Dense environments such as local galaxy clusters tend to host larger populations of quiescent galaxies than the field (e.g. Dressler 1980; Poggianti et al. 1999; Balogh et al. 2004, 2009, 2016; Kawinwanichakij et al. 2017; Pintos-Castro et al. 2019; Ragusa et al. 2025). At the fraction of quenched galaxies in clusters decreases with increasing redshift (e.g. Balogh et al. 2004; Raichoor and Andreon 2012; Pintos-Castro et al. 2019) but always remains above the field level. The situation is less clear at . Some authors find that galaxy clusters already exhibit elevated levels of quenching (e.g. Newman et al. 2014; Cooke et al. 2019; Lemaux et al. 2019; Strazzullo et al. 2019; van der Burg et al. 2020; Toni et al. 2025), while others observe no difference with field levels before (e.g. Tran et al. 2010; Brodwin et al. 2013; Alberts et al. 2016; Nantais et al. 2020; Trudeau et al. 2024).
@@
Another example of the interplay between galaxy evolution and cluster environment is the Brightest Cluster Galaxy (BCG). BCGs are the most massive galaxies in the Universe and they are found close to the centre of the potential wells of galaxy clusters (Zitrin et al. 2012; Hashimoto et al. 2014; Cui et al. 2016; Lopes et al. 2018). Their properties (e.g. shape, mass, star-formation history) tend to correlate with those of their host clusters (e.g. Ebeling et al. 2021). A particular example is the relation between AGN activity in BCGs and the regulation of the temperature of the intracluster medium (ICM, Hu et al. 1985; Burns 1990; Cavagnolo et al. 2008; Rafferty et al. 2008; Hlavacek-Larrondo et al. 2012). An imbalance between the cooling of the ICM and the energy injected by the AGN (e.g. McNamara and O’Connell 1989; Fabian 1994; McNamara et al. 2005; Rafferty et al. 2006) can trigger gas condensation onto the BCG, a phenomenon called a ‘cooling flow.’
@@
BCG evolution is intimately linked to the evolution of their host clusters. At , BCG evolution is dominated by gas-poor mergers (e.g. Aragón-Salamanca et al. 1998; Dubinski 1998; Lidman et al. 2012, 2013; Burke and Collins 2013; Golden-Marx and Miller 2018; Golden-Marx et al. 2025; Montenegro-Taborda et al. 2023). While most local BCGs are quiescent, between 20 and 35% of them exhibit small amounts of star formation (usually below the field level, see Orellana-González et al. 2022) powered by residual cooling flows (e.g. Crawford et al. 1999; Rawle et al. 2012; Oliva-Altamirano et al. 2014; McDonald et al. 2016).
@@
However, the formation and early evolution of BCGs remain poorly understood. A popular formation model (De Lucia and Blaizot 2007) predicted that 80% of the stellar mass of BCGs would form before in separate progenitors that progressively assemble into a BCG through gas-poor mergers. Observations of BCGs showed a different picture: high-redshift BCGs are diverse, but some of them display substantial levels of in-situ star formation (Webb et al. 2015a; McDonald et al. 2016; Bonaventura et al. 2017). The cause of the star formation is not currently understood. Some authors (e.g. Rennehan et al. 2020; Remus et al. 2023) suggest that it might be powered by the initial formation of the BCG in a gas-rich merger involving multiple galaxies. Observations of protocluster cores with elevated star-formation rates (SFRs) and interacting galaxies support this scenario (Miley et al. 2006; Kuiper et al. 2011; Miller et al. 2018; Coogan et al. 2023). One cluster core at , SpARCS104922.6+564032.5 (Webb et al. 2015b; Hlavacek-Larrondo et al. 2020) indicates that massive cooling flows could also power star formation.
@@
Our current understandings of quenching in high-redshift clusters and BCG early evolution are limited by small numbers statistics since most cluster samples contain only a few objects at (e.g. Wilson et al. 2006, 2009; Wylezalek et al. 2013; Bleem et al. 2015; Pierre et al. 2016; Bulbul et al. 2022, 2024). There are several recent protocluster searches that explore these redshifts (e.g. Martinache et al. 2018; Ouchi et al. 2018; Toshikawa et al. 2018; Golden-Marx et al. 2019; Ando et al. 2020; Gully et al. 2024) but their sample purities tend to be low (e.g. Hung et al. 2025).
@@
With the advent of a new generation of large-scale survey facilities (Euclid, Nancy Grace Roman, Vera C. Rubin, see Spergel et al. 2015; Ivezić et al. 2019; Robertson et al. 2019; Euclid Collaboration: Scaramella et al. 2022; Grishin et al. 2025), this situation is changing and large samples of (proto)clusters are becoming available. Euclid is the only one of these telescopes that is currently in service. Located at the second Sun-Earth Lagrange point, Euclid has a 1.2-metre primary mirror (Euclid Collaboration: Mellier et al. 2025). By the end of its six-year mission, Euclid is expected to observe deg2 of the sky (Euclid Collaboration: Mellier et al. 2025) and to find about galaxy clusters at according to the latest estimates (Euclid Collaboration: Mellier et al., 2025; see also the estimations of Euclid Collaboration: Adam et al. 2019 and Sartoris et al. 2016). The Euclid search for galaxy clusters uses two algorithms: the Adaptive Matched Identifier of Clustered Objects (AMICO) and PZWav (Euclid Collaboration: Adam et al., 2019; Euclid Collaboration: Bhargava et al., 2025). AMICO is a matched filter algorithm: it detects clusters with a redshift-dependent filter mimicking the sum of the cluster contribution and the noise (Bellagamba et al., 2018; Maturi et al., 2019). PZWav applies instead a spatial filter constructed from the difference of two Gaussians. The width of the Gaussians are chosen to select overdensities with sizes consistent with a cluster extent on the sky (Thongkham et al., 2024b).
@@
PZWav is also the algorithm used by another large-scale cluster survey, the Massive and Distant Clusters of WISE Survey 2 (MaDCoWS2; Thongkham et al., 2024b, a). MaDCoWS2 derives photometric redshifts (Brodwin et al., 2006) using data from the Dark Energy Camera (DECam) and the Wide-Field Infrared Survey Explorer (WISE), and uses these photometric redshifts as input to PZWav. While the MaDCoWS2 survey identified 6959 candidate clusters at , WISE resolution (PSF FWHM of and in the two bluest channels, see Wright et al., 2010) is insufficient to distinguish individual galaxies in the packed cores of high-redshift clusters. Near-infrared data sets with better resolution, such as Euclid images, are needed to enable galaxy evolution studies of MaDCoWS2 clusters. Furthermore, the Euclid Wide Survey will largely overlap with the MaDCoWS2 footprint (Euclid Collaboration: Scaramella et al., 2022; Euclid Collaboration: Mellier et al., 2025).
@@
In this paper, we present Euclid’s first view of a candidate cluster both detected in MaDCoWS2 and Euclid preliminary cluster search: MOO2 J0337428386/EUCLQ1CL J033730.18283827.6, which we nickname the ‘Puddle’ cluster due to the intriguing BCG complex revealed by Euclid. The left panel of Fig. 1 presents a tricolour image based on Euclid data; the right panel shows a zoomed, view of the BCG complex. This system represents a test of the capabilities – and limitations – of Euclid for the discovery and follow-up of high-redshift galaxy clusters.
@@
The paper is divided as follows: Sect. 2 presents the data used to discover and characterise the Puddle cluster. Section 3 presents the data analysis, focusing on the galaxy population and the central structure (i.e. the BCG complex) characterisation. Section 4 discusses our findings, which are summarised in Sect. 5. Throughout this paper, we assume Planck Collaboration: Aghanim et al. (2020) CDM cosmology as implemented in astropy.cosmology: and . At , an angular distance of corresponds to a proper distance between 1.03 and 1.04 Mpc. All magnitudes are in the AB system. We assume a Kroupa and Boily (2002) initial mass function.
@@ 2.1 MaDCoWS2 catalogue
MaDCoWS2 is an optical and near-infrared catalogue that covers and contains 133 036 candidate galaxy clusters (Thongkham et al., 2024b, a). It is based on the bands from the DECam Legacy Survey (DECaLS; see Flaugher et al., 2015; Dey et al., 2019) and the bands from the CatWISE2020 catalogue (Eisenhardt et al., 2020; Marocco et al., 2021). The photometric redshift probability distribution functions (PDFs; see Brodwin et al., 2006) computed from these bands are input to PZWav. In MaDCoWS2, the Puddle cluster is detected as a overdensity with a signal-to-noise ratio (S/N) of 5.6.
@@ 2.2 Euclid Q1 data release
Euclid’s two instruments are the Visible Camera (VIS, Euclid Collaboration: Cropper et al., 2025), a single-band (denoted ) imaging camera covering the to Å range, and the Near-Infrared Spectrometer and Photometer (NISP, Euclid Collaboration: Jahnke et al., 2025). NISP can perform imaging (, , and bands) and multiobject slitless spectroscopy (blue and red grisms).
The Euclid Consortium Quick Release 1 (hereafter the Q1 release, Euclid Quick Release Q1, 2025; Euclid Collaboration: Mellier et al., 2025; Euclid Collaboration: Aussel et al., 2025) is composed primarily of Euclid images, photometric and spectroscopic catalogues (MER and SPE catalogues, see Euclid Collaboration: Aussel et al., 2025; Euclid Collaboration: Romelli et al., 2025; Euclid Collaboration: Copin et al., 2025) of the Euclid Deep Field North, Euclid Deep Field Fornax, and Euclid Deep Field South. The two latter fields are also covered by the MaDCoWS2 catalogue. The Puddle cluster is located in the Euclid Deep Field Fornax.
A preliminary cluster search is also part of the Q1 release (Euclid Collaboration: Bhargava et al., 2025). This search focuses on detections. Euclid photometric redshifts are still being refined and currently have large uncertainties at . However, Euclid Collaboration: Bhargava et al. (2025) includes a sample of 15 unconfirmed clusters that meet the requirement of being detected both by PZWav and AMICO. The Puddle cluster is among this list. It is detected by PZWav as a overdensity with an S/N of 9.17 and by AMICO as a overdensity with an S/N of 16.39 and a richness of . AMICO and PZWAv compute S/N differently; hence AMICO produces systematically higher S/N than PZWAv (see Fig. 4 of Euclid Collaboration: Bhargava et al. 2025). However, no redshift uncertainties are provided, and the scatter between photometric and spectroscopic redshifts is estimated for only. We therefore use the MaDCoWS2 detection as our primary reference for the cluster position and photometric redshift.
@@ 2.2.1 MER catalogue
The Q1 MER catalogue111available in Euclid Science Archive System, at https://eas.esac.esa.int/sas/ combines VIS, NISP, and several external data sets, including bands from DECam (Honscheid and DePoy, 2008; Flaugher et al., 2015) on the 4 m Blanco Telescope, to create a common catalogue with consistently measured photometry. Euclid and external images are organised in calibrated tiles with overlaps and pixel scale (Euclid Collaboration: Romelli et al., 2025). Sources are detected in VIS and NISP, de-blended, and organised into a single source list which is then used to measure photometry in all bands.
To identify the presence of a red sequence, we use the and bands, which bracket the Å break at . While the band lies blueward of the Å break at these redshifts, its broadness (effective width of Å, with a mean wavelength of Å, see Rodrigo et al., 2012; Rodrigo and Solano, 2020; Rodrigo et al., 2024) limits its usefulness for characterising galaxy populations. Specifically, we use the and fluxes obtained via a single Sérsic model-fitting, denoted by a ‘_sersic’ suffix in the MER catalogue.

### 13 Mar 2026 v3

@@
We present an example cluster follow-up study with Euclid. Our target, a candidate cluster nicknamed the ‘Puddle’, was initially discovered by the Massive and Distant Clusters of WISE Survey 2 as a candidate cluster. It was also detected independently as a candidate with the two cluster-finding algorithms in Euclid Quick Release 1 (Q1). A Keck MOSFIRE spectrum shows the brightest nucleus is at and is dominated by an active galactic nucleus. Our analysis focused on the galaxy population and the brightest cluster galaxy (BCG), and is based on Euclid and ancillary photometry. Compared to similar fields, we measured an overdensity of galaxies with in a radius around the BCG. About % of the completeness-corrected galaxy population is red, which is consistent with some clusters at but lower than others. Euclid imaging revealed that six or seven galaxies appear to be assembling to form the future BCG. Spectral energy distribution fitting suggests that the merging BCG has a stellar mass of and that it experienced a short burst of star formation Myr ago. Its morphology and star-formation history suggest that the proto-BCG is a more evolved version of the merging core of SPT234956. These systems indicate that multiobject mergers might be a common BCG formation process. Assuming a similar density of mergers in the Euclid Wide Survey, we expect that Euclid will discover approximately 400 assembling BCGs by the end of its mission.
@@
Galaxy clusters sit at the nodes of the cosmic web and are linked by filaments (Bond et al. 1996; Springel et al. 2005; Vogelsberger et al. 2014). As the maxima of the density field in the Universe, galaxy clusters are natural laboratories to test a variety of astrophysical phenomena in different fields: cosmology (e.g. Vikhlinin et al. 2009; Allen et al. 2011; Pierre et al. 2016), plasma and high-energy physics (e.g. McNamara and Nulsen 2007, 2012; Zhuravleva et al. 2014), and galaxy evolution (e.g. von der Linden et al. 2010; Peng et al. 2010b, 2012).
@@
The cessation of star formation in galaxies, usually called ‘quenching’, depends on two main factors: galaxy mass (Kauffmann et al. 2003) and environment density (Peng et al. 2010b, 2012). Dense environments such as local galaxy clusters tend to host a larger fraction of quiescent galaxies than field galaxy samples (e.g. Dressler 1980; Poggianti et al. 1999; Balogh et al. 2004, 2009, 2016; Kawinwanichakij et al. 2017; Pintos-Castro et al. 2019; Ragusa et al. 2025). At the fraction of quenched galaxies in clusters decreases with increasing redshift (e.g. Balogh et al. 2004; Raichoor and Andreon 2012; Pintos-Castro et al. 2019) but always remains above the field level. The situation is less clear at . Some authors find that high-redshift galaxy clusters already exhibit elevated levels of quenching (e.g. Newman et al. 2014; Cooke et al. 2019; Lemaux et al. 2019; Strazzullo et al. 2019; van der Burg et al. 2020; Toni et al. 2026), while others observe no difference with field levels before (e.g. Tran et al. 2010; Brodwin et al. 2013; Alberts et al. 2016; Nantais et al. 2020; Trudeau et al. 2024).
@@
Brightest cluster galaxies (BCGs) represent another example of the interplay between galaxy evolution and cluster environment. BCGs are the most massive galaxies in the Universe, and they are found close to the centre of the potential wells of galaxy clusters (Zitrin et al. 2012; Hashimoto et al. 2014; Cui et al. 2016; Lopes et al. 2018). Their properties (e.g. shape, mass, star-formation history) tend to correlate with those of their host clusters (e.g. Ebeling et al. 2021). A particular example is the relation between the active galactic nucleus (AGN) activity in BCGs and the regulation of the temperature of the intracluster medium (ICM; Hu et al. 1985; Burns 1990; Cavagnolo et al. 2008; Rafferty et al. 2008; Hlavacek-Larrondo et al. 2012). An imbalance between the cooling of the ICM and the energy injected by the AGN (e.g. McNamara and O’Connell 1989; Fabian 1994; McNamara et al. 2005; Rafferty et al. 2006) can trigger gas condensation onto the BCG, a phenomenon called a ‘cooling flow’.
@@
The BCG evolution is intimately linked to the evolution of its host cluster. At , BCG evolution is dominated by gas-poor mergers (e.g. Aragón-Salamanca et al. 1998; Dubinski 1998; Lidman et al. 2012, 2013; Burke and Collins 2013; Golden-Marx and Miller 2018; Golden-Marx et al. 2025; Montenegro-Taborda et al. 2023). While most local BCGs are quiescent, between 20% and 35% of them exhibit small amounts of star formation (usually below the field level, see Orellana-González et al. 2022) powered by residual cooling flows (e.g. Crawford et al. 1999; Rawle et al. 2012; Oliva-Altamirano et al. 2014; McDonald et al. 2016).
@@
However, the formation and early evolution of BCGs remain poorly understood. A popular formation model (De Lucia and Blaizot 2007) predicted that 80% of the stellar mass of BCGs would form before in separate progenitors that progressively assemble into a BCG through gas-poor mergers. Observations of BCGs have shown a different picture: High-redshift BCGs are diverse, but some of them display substantial levels of in situ star formation (Webb et al. 2015a; McDonald et al. 2016; Bonaventura et al. 2017). The cause of the star formation is not currently understood. Some authors (e.g. Rennehan et al. 2020; Remus et al. 2023) suggest that it might be powered by the initial formation of the BCG in a gas-rich merger involving multiple galaxies. Observations of protocluster cores with elevated star-formation rates (SFRs) and interacting galaxies support this scenario (Miley et al. 2006; Kuiper et al. 2011; Miller et al. 2018; Coogan et al. 2023). One cluster core at , SpARCS104922.6+564032.5 (Webb et al. 2015b; Hlavacek-Larrondo et al. 2020) indicates that massive cooling flows could also power star formation.
@@
Our current understanding of quenching in high-redshift clusters and BCG early evolution are limited by small number statistics since most cluster samples contain only a few objects at (e.g. Wilson et al. 2006, 2009; Wylezalek et al. 2013; Bleem et al. 2015; Pierre et al. 2016; Bulbul et al. 2022, 2024). There are several recent protocluster searches that explore these redshifts (e.g. Martinache et al. 2018; Ouchi et al. 2018; Toshikawa et al. 2018; Golden-Marx et al. 2019; Ando et al. 2020; Gully et al. 2024), but their sample purities tend to be low (e.g. Hung et al. 2025).
@@
With the advent of a new generation of large-scale survey facilities (Euclid, Nancy Grace Roman, Vera C. Rubin, see Spergel et al. 2015; Ivezić et al. 2019; Robertson et al. 2019; Euclid Collaboration: Scaramella et al. 2022; Grishin et al. 2025), this situation is changing, and large samples of (proto)clusters are becoming available. Euclid is the only one of these facilities that is currently in service. Located at the second Sun-Earth Lagrange point, Euclid has a 1.2-metre primary mirror (Euclid Collaboration: Mellier et al. 2025). By the end of its six-year mission, Euclid is expected to observe deg2 of the sky (Euclid Collaboration: Mellier et al. 2025) and to find about galaxy clusters at , according to the latest estimates (Euclid Collaboration: Mellier et al. 2025, see also the estimations of Euclid Collaboration: Adam et al. 2019 and Sartoris et al. 2016).
The Euclid search for galaxy clusters uses two algorithms: the Adaptive Matched Identifier of Clustered Objects (AMICO) and PZWav (Euclid Collaboration: Adam et al. 2019; Euclid Collaboration: Bhargava et al. 2025). AMICO is a matched filter algorithm and it detects clusters with a redshift-dependent filter mimicking the sum of the cluster contribution and the noise (Bellagamba et al. 2018; Maturi et al. 2019). PZWav instead applies a spatial filter constructed from the difference of two Gaussians. The width of the Gaussians was chosen to select overdensities with sizes consistent with a cluster extent on the sky (Thongkham et al. 2024b).
@@
The PZWav algorithm is also used by another large-scale cluster survey, the Massive and Distant Clusters of WISE Survey 2 (MaDCoWS2; Thongkham et al. 2024b, a). MaDCoWS2 derives photometric redshifts (Brodwin et al. 2006) using data from the Dark Energy Camera (DECam) and the Wide-Field Infrared Survey Explorer (WISE), and it uses these photometric redshifts as input to PZWav. While the MaDCoWS2 survey identified 6959 candidate clusters at , the resolution of WISE – the point spread functions (PSFs) full widths at half maxima (FWHM) are and in the two bluest channels (see Wright et al. 2010) – is insufficient for distinguishing individual galaxies in the packed cores of high-redshift clusters. Near-infrared data sets with better resolution, such as Euclid images, are needed to enable galaxy evolution studies of MaDCoWS2 clusters. Furthermore, the Euclid Wide Survey will largely overlap with the MaDCoWS2 footprint (Euclid Collaboration: Scaramella et al. 2022; Euclid Collaboration: Mellier et al. 2025).
@@
In this paper, we present Euclid’s first view of a candidate cluster detected in both MaDCoWS2 and the Euclid preliminary cluster search: MOO2 J0337428386/EUCLQ1CL J033730.18283827.6. We nicknamed this candidate the ‘Puddle’ cluster due to the intriguing BCG complex revealed by Euclid. The left panel of Fig. 1 presents a tricolour image based on Euclid data; the right panel shows a zoomed view of the BCG complex. This system represents a test of the capabilities – and limitations – of Euclid for the discovery and follow-up of high-redshift galaxy clusters.
@@
The paper is divided as follows: Sect. 2 presents the data used to discover and characterise the Puddle cluster. Section 3 presents the data analysis, focusing on the galaxy population and the central structure (i.e. the BCG complex) characterisation. In Sect. 4 we discuss our findings, which are summarised in Sect. 5. Throughout this paper, we assume a Planck Collaboration: Aghanim et al. (2020) cold dark matter (CDM) cosmology as implemented in astropy.cosmology: and . At , an angular distance of corresponds to a proper distance between 1.03 and 1.04 Mpc. All magnitudes are in the AB system. We assumed a Kroupa and Boily (2002) initial mass function.
@@ 2.1 MaDCoWS2 catalogue
MaDCoWS2 is an optical and near-infrared catalogue that covers and contains 133 036 candidate galaxy clusters (Thongkham et al. 2024b, a). It is based on the bands from the DECam Legacy Survey (DECaLS; see Flaugher et al. 2015; Dey et al. 2019) and the bands from the CatWISE2020 catalogue (Eisenhardt et al. 2020; Marocco et al. 2021). The photometric redshift probability distribution functions (PDFs; see Brodwin et al. 2006) computed from these bands was input in PZWav. In MaDCoWS2, the Puddle cluster was detected as a overdensity with a signal-to-noise ratio (S/N) of 5.6.
@@ 2.2 Euclid Q1 data release
Euclid’s two instruments are the Visible Camera (VIS; Euclid Collaboration: Cropper et al. 2025), a single-band (denoted ) imaging camera covering the to Å range, and the Near-Infrared Spectrometer and Photometer (NISP; Euclid Collaboration: Jahnke et al. 2025). The NISP can perform imaging (, , and bands) and multiobject slitless spectroscopy (blue and red grisms).
The Euclid Consortium Quick Release 1 (hereafter the Q1 release, Euclid Quick Release Q1 2025; Euclid Collaboration: Mellier et al. 2025; Euclid Collaboration: Aussel et al. 2025) is composed primarily of Euclid images and photometric and spectroscopic catalogues (MER and SPE catalogues, see Euclid Collaboration: Aussel et al. 2025; Euclid Collaboration: Romelli et al. 2025; Euclid Collaboration: Copin et al. 2025) of the Euclid Deep Field North, Euclid Deep Field Fornax, and Euclid Deep Field South. The two last fields are also covered by the MaDCoWS2 catalogue. The Puddle cluster is located in the Euclid Deep Field Fornax.
A preliminary cluster search is also part of the Q1 release (Euclid Collaboration: Bhargava et al. 2025). This search focuses on detections. Euclid photometric redshifts are still being refined and currently have large uncertainties at . However, Euclid Collaboration: Bhargava et al. (2025) includes a sample of 15 unconfirmed clusters that meet the requirement of being detected both by PZWav and AMICO. The Puddle cluster is among this list. It was detected by PZWav as a overdensity with an S/N of 9.17 and by AMICO as a overdensity with an S/N of 16.39 and a richness of . AMICO and PZWAv compute S/N differently; hence AMICO produces a systematically higher S/N than PZWAv (see Fig. 4 of Euclid Collaboration: Bhargava et al. 2025). However, no redshift uncertainties are provided, and the scatter between photometric and spectroscopic redshifts is estimated for only. We therefore used the MaDCoWS2 detection as our primary reference for the cluster position and photometric redshift.
@@ 2.2.1 MER catalogue
The Q1 MER catalogue111available in Euclid Science Archive System, at https://eas.esac.esa.int/sas/ combines VIS, NISP, and several external data sets, including bands from DECam (Honscheid and DePoy 2008; Flaugher et al. 2015) on the 4 m Blanco Telescope, to create a common catalogue with consistently measured photometry. Euclid and external images are organised in calibrated tiles with overlaps and pixel scale (Euclid Collaboration: Romelli et al. 2025). Sources are detected in VIS and NISP, de-blended, and organised into a single source list which is then used to measure photometry in all bands.
To identify the presence of a red sequence, we used the and bands, which bracket the Å break at . While the band lies blueward of the Å break at these redshifts, its broadness (effective width of Å, with a mean wavelength of Å, see Rodrigo et al. 2012; Rodrigo and Solano 2020; Rodrigo et al. 2024) limits its usefulness for characterising galaxy populations. Specifically, we used the and fluxes obtained via a single Sérsic model-fitting, denoted by a ‘_sersic’ suffix in the MER catalogue.